# Stationary and Non-Stationary Transition Probabilities in Decision Making: Modeling COVID-19 Dynamics

### v1

@@ 3.3 Stationary Transition Probabilities
  5. (e)
    The probability of landing in State $R$ after leaving State $I_{s}$ is as follows:
    $\mathbb{P}(R|I_{s})=\displaystyle\frac{\mathbb{P}(R\cap I_{s})}{\mathbb{P}(I_{s})},$
    where $\mathbb{P}(I_{s})=\mathbb{P}({\mathcal{V}}_{\delta_{5},\delta_{6},\delta_{7}})$ and $\mathbb{P}(R\cap I_{s})=\begin{pmatrix}\delta_{5}+\delta_{6}+\delta_{7}\\ ~{}\\ \delta_{5}\end{pmatrix}\tau_{is,r}^{\delta_{5}}(1-\tau_{is,r})^{\delta_{6}+\delta_{7}}$.
    Therefore, the transition probability from state $I_{s}$ to state $R$ is given by (3.18).
  6. (f)
    The probability of landing in State $H$ after leaving State $I_{s}$ is as follows:
    $\mathbb{P}(H|I_{s})=\displaystyle\frac{\mathbb{P}(H\cap I_{s})}{\mathbb{P}(I_{s})},$
    where $\mathbb{P}(I_{s})=\mathbb{P}({\mathcal{V}}_{\delta_{5},\delta_{6},\delta_{7}})$ and $\mathbb{P}(H\cap I_{s})=\begin{pmatrix}\delta_{5}+\delta_{6}+\delta_{7}\\ ~{}\\ \delta_{6}\end{pmatrix}\tau_{is,h}^{\delta_{6}}(1-\tau_{is,h})^{\delta_{5}+\delta_{7}}$.
    Therefore, the transition probability from state $I_{s}$ to state $H$ is given by (3.19).
  7. (g)
    The probability of landing in State $D$ after leaving State $I_{s}$ is as follows:
    $\mathbb{P}(D|I_{s})=\displaystyle\frac{\mathbb{P}(D\cap I_{s})}{\mathbb{P}(I_{s})},$
    where $\mathbb{P}(I_{s})=\mathbb{P}({\mathcal{V}}_{\delta_{5},\delta_{6},\delta_{7}})$ and $\mathbb{P}(D\cap I_{s})=\begin{pmatrix}\delta_{5}+\delta_{6}+\delta_{7}\\ ~{}\\ \delta_{7}\end{pmatrix}\tau_{is,d}^{\delta_{7}}(1-\tau_{is,d})^{\delta_{5}+\delta_{6}}$.
    Therefore, the transition probability from state $I_{s}$ to state $D$ is given by (3.20).
  8. (h)
    The probability of landing in State $D$ after leaving State $H$ is as follows:
    $\mathbb{P}(D|H)=\displaystyle\frac{\mathbb{P}(D\cap H)}{\mathbb{P}(H)},$
    where $\mathbb{P}(H)=\mathbb{P}({\mathcal{V}}_{\delta_{8},\delta_{9}})$ and $\mathbb{P}(D\cap H)=\begin{pmatrix}\delta_{8}+\delta_{9}\\ ~{}\\ \delta_{8}\end{pmatrix}\tau_{h,d}^{\delta_{8}}(1-\tau_{h,d})^{\delta_{9}}$.
    Therefore, the transition probability from state $H$ to state $D$ is given by (3.21).
  9. (i)
    Transition Probability From State $H$ to State $R$
    The probability of landing in State $R$ after leaving State $H$ is as follows:
    $\mathbb{P}(R|H)=\displaystyle\frac{\mathbb{P}(R\cap H)}{\mathbb{P}(H)},$
    where $\mathbb{P}(H)=\mathbb{P}({\mathcal{V}}_{\delta_{8},\delta_{9}})$ and $\mathbb{P}(R\cap H)=\begin{pmatrix}\delta_{8}+\delta_{9}\\ ~{}\\ \delta_{9}\end{pmatrix}\tau_{h,r}^{\delta_{9}}(1-\tau_{h,r})^{\delta_{8}}$.
    Therefore, the transition probability from state $H$ to state $R$ is given by (3.22).
  10. (j)
@@
Stationary transition probabilities also assume that the process reaches a stable equilibrium where the disease dynamics are predictable. This model is not a good fit for predicting the long term dynamics of Covid-19 due to ongoing uncertainties such as: the evolution of the virus with new variants and changes in viral behavior; the public health policies with implementation, adjustment and measures continuously influence disease dynamics; the dynamics of immunity, including the duration of vaccine-induced immunity and natural immunity, are evolving, affecting disease spread and transition rates. Therefore, the need of non-stationary transition probabilities for our study.
@@ 3.4 Non-stationary Transition Probabilities
In contrast to stationary transition probabilities, non-stationary transition probabilities change over time. This means that the likelihood of moving from one state to another can vary due to factors such as time, evolving disease dynamics, intervention strategies, and seasonal effects. In this study, we focus on time as the primary factor influencing our non-stationary transition probabilities. Specifically, we model the various probabilistic parameters as functions of time to achieve a more accurate and realistic representation of disease dynamics and the impact of interventions.
@@
    $\begin{array}[]{rllll}\mathbb{P}(E|S,t)&=&\begin{pmatrix}|S|_{k}\\ ~{}\\ \delta_{1}\end{pmatrix}\left(\displaystyle\frac{(\zeta_{ia,s}(t)+\zeta_{is,s}(t)+\zeta_{h,s}(t))(|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k})}{N-|D|_{k}}\right)^{\delta_{1}}\\ {}\hfil\\ &&\left(1-\displaystyle\frac{(\zeta_{ia,s}(t)+\zeta_{is,s}(t)+\zeta_{h,s}(t))(|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k})}{N-|D|_{k}}\right)^{|S|_{k}-\delta_{1}},\end{array}$
@@
    $\mathbb{P}(I_{a}|E)=\begin{pmatrix}|E|_{k}\\ ~{}\\ \delta_{2}\end{pmatrix}\tau_{e,ia}^{\delta_{2}}(t)(1-\tau_{e,ia}(t))^{|E|_{k}-\delta_{2}},$
@@
    $\mathbb{P}(I_{s}|I_{a})=\displaystyle\frac{\begin{pmatrix}\delta_{3}+\delta_{4}\\ ~{}\\ \delta_{3}\end{pmatrix}\tau_{ia,is}(t)^{\delta_{3}}(1-\tau_{ia,is}(t))^{\delta_{4}}}{{\mathcal{M}}_{|I_{a}|_{k},\delta_{3},\delta_{4}}\tau_{ia,is}(t)^{\delta_{3}}\tau_{ia,r}(t)^{\delta_{4}}(1-(\tau_{ia,is}(t)+\tau_{ia,r}(t)))^{|I_{a}|_{k}-(\delta_{3}+\delta_{4})}},$
@@
    $\mathbb{P}(R|I_{a})=\displaystyle\frac{\begin{pmatrix}\delta_{3}+\delta_{4}\\ ~{}\\ \delta_{4}\end{pmatrix}\tau_{ia,r}(t)^{\delta_{4}}(1-\tau_{ia,r}(t))^{\delta_{3}}}{{\mathcal{M}}_{|I_{a}|_{k},\delta_{3},\delta_{4}}\tau_{ia,is}(t)^{\delta_{3}}\tau_{ia,r}(t)^{\delta_{4}}(1-(\tau_{ia,is}(t)+\tau_{ia,r}(t)))^{|I_{a}|_{k}-(\delta_{3}+\delta_{4})}},$
@@
    $\mathbb{P}(R|I_{s})=\displaystyle\frac{\begin{pmatrix}\delta_{5}+\delta_{6}+\delta_{7}\\ ~{}\\ \delta_{5}\end{pmatrix}\tau_{is,r}(t)^{\delta_{5}}(1-\tau_{is,r}(t))^{\delta_{6}+\delta_{7}}}{{\mathcal{M}}_{|I_{s}|_{k},\delta_{5},\delta_{6},\delta_{7}}\tau_{is,r}(t)^{\delta_{5}}\tau_{is,h}(t)^{\delta_{6}}\tau_{is,d}(t)^{\delta_{7}}(1-(\tau_{is,r}(t)+\tau_{is,h}(t)+\tau_{is,d}(t)))^{|I_{s}|_{k}-(\delta_{5}+\delta_{6}+\delta_{7})}},$
@@
    $\mathbb{P}(H|I_{s})=\displaystyle\frac{\begin{pmatrix}\delta_{5}+\delta_{6}+\delta_{7}\\ ~{}\\ \delta_{6}\end{pmatrix}\tau_{is,h}(t)^{\delta_{6}}(1-\tau_{is,h}(t))^{\delta_{5}+\delta_{7}}}{{\mathcal{M}}_{|I_{s}|_{k},\delta_{5},\delta_{6},\delta_{7}}\tau_{is,r}(t)^{\delta_{5}}\tau_{is,h}(t)^{\delta_{6}}\tau_{is,d}(t)^{\delta_{7}}(1-(\tau_{is,r}(t)+\tau_{is,h}(t)+\tau_{is,d}(t)))^{|I_{s}|_{k}-(\delta_{5}+\delta_{6}+\delta_{7})}},$
@@
    $\mathbb{P}(D|I_{s})=\displaystyle\frac{\begin{pmatrix}\delta_{5}+\delta_{6}+\delta_{7}\\ ~{}\\ \delta_{7}\end{pmatrix}\tau_{is,d}(t)^{\delta_{7}}(1-\tau_{is,d}(t))^{\delta_{5}+\delta_{6}}}{{\mathcal{M}}_{|I_{s}|_{k},\delta_{5},\delta_{6},\delta_{7}}\tau_{is,r}(t)^{\delta_{5}}\tau_{is,h}(t)^{\delta_{6}}\tau_{is,d}(t)^{\delta_{7}}(1-(\tau_{is,r}(t)+\tau_{is,h}(t)+\tau_{is,d}(t)))^{|I_{s}|_{k}-(\delta_{5}+\delta_{6}+\delta_{7})}},$
@@
    $\mathbb{P}(D|H)=\displaystyle\frac{\begin{pmatrix}\delta_{8}+\delta_{9}\\ ~{}\\ \delta_{8}\end{pmatrix}\tau_{h,d}(t)^{\delta_{8}}(1-\tau_{h,d}(t))^{\delta_{9}}}{{\mathcal{M}}_{|H|_{k},\delta_{8},\delta_{9}}\tau_{h,d}(t)^{\delta_{8}}\tau_{h,r}(t)^{\delta_{9}}(1-(\tau_{h,d}(t)+\tau_{h,r}(t)))^{|H|_{k}-(\delta_{8}+\delta_{9})}},$
@@
    $\mathbb{P}(R|H)=\displaystyle\frac{\begin{pmatrix}\delta_{8}+\delta_{9}\\ ~{}\\ \delta_{9}\end{pmatrix}\tau_{h,r}(t)^{\delta_{9}}(1-\tau_{h,r}(t))^{\delta_{8}}}{{\mathcal{M}}_{|H|_{k},\delta_{8},\delta_{9}}\tau_{h,d}(t)^{\delta_{8}}\tau_{h,r}(t)^{\delta_{9}}(1-(\tau_{h,d}(t)+\tau_{h,r}(t)))^{|H|_{k}-(\delta_{8}+\delta_{9})}},$
@@ 4 COVID-19 Dynamics Over 365 Days
    Figure 3: Low Values of $t$ (from 00 up to $2.18$)
@@
    Figure 4: Moderate to High Values of $t$ ($2.18$ and above)
    When $t$ reaches $2.18$ and beyond, the dynamics settle. The susceptible population remains consistently larger than the deceased population, due to a slower increase and decrease in the deceased and susceptible compartments, respectively. This phase signifies a balance between disease spread and recovery, resulting in a more controlled epidemic progression. By approximately day 250, the system reaches a steady state, with variations in all compartments approaching zero. Both the susceptible and deceased populations balance, indicating that the epidemic has subsided and the populations are no longer undergoing significant changes.
These plots illustrate the progression of individuals across different compartments of a COVID-19 model over time (in days), offering valuable insights for understanding and predicting disease dynamics, and for planning public health interventions. Overall, as the probabilistic function increases (when $t$ is in the range $[0,0.4)$), there is a rapid decline in the susceptible population, reflecting a high rate of new infections, while the deceased population rises quickly, indicating a severe impact of the disease and a high transition rate from infection to death. This rapid decline in susceptibles and fast increase in deceased population reduce the number of active infections and limit the size of the recovered population, thereby influencing the overall dynamics. When the probabilistic function decreases (for $t$ in the range $[0.4,4.2)$), the decline in the susceptible population slows, indicating a reduction in the infection rate, and the increase in the deceased population also slows, reflecting a less aggressive disease progression. This leads to a more balanced dynamic, with intermediate compartments like the infected population maintaining a more consistent size. Finally, when the probabilistic function becomes constant (approaching $0.15$ for $t$ in the range $[4.2,\infty)$), the sizes of the susceptible and deceased populations settle, indicating that the system has reached an equilibrium with no significant new infections or deaths occurring. This constancy suggests that other compartments, such as the infected and recovered populations, have also leveled off, leading to a long-term steady state in the epidemic’s dynamics.

### v2

@@ 3.3 Stationary Transition Probabilities
Remark 3.2
The transition probabilities (3.18)-(3.22) are obtained following the same reasoning as in Part (c) and (d).
Remark 3.3
@@
Stationary transition probabilities assume that the process reaches a stable equilibrium where the disease dynamics are predictable. This model is not a good fit for predicting the long term dynamics of Covid-19 due to ongoing uncertainties such as: the evolution of the virus with new variants and changes in viral behavior; the public health policies with implementation, adjustment and measures continuously influence disease dynamics; the dynamics of immunity, including the duration of vaccine-induced immunity and natural immunity, are evolving, affecting disease spread and transition rates. Therefore, the need of non-stationary transition probabilities for our study.
@@ 3.4 Non-stationary Transition Probabilities
In contrast to stationary transition probabilities, non-stationary transition probabilities change over time. This means that the likelihood of moving from one state to another can vary due to factors such as time, evolving disease dynamics, intervention strategies, and seasonal effects. In this study, we focus on time as the primary factor influencing our non-stationary transition probabilities. Specifically, we model the various probabilistic parameters as functions of time to achieve a more accurate and realistic representation of disease dynamics and the impact of interventions. In what follows,
$\beta(t)=1-\tau_{is,r}(t)-\tau_{is,h}(t)-\tau_{is,d}(t).$
@@
    $\begin{array}[]{rllll}\mathbb{P}(E|S,t)&=&\begin{pmatrix}|S|_{k}\\ ~\\ \delta_{1}\end{pmatrix}\left(\displaystyle\frac{(\zeta_{ia,s}(t)+\zeta_{is,s}(t)+\zeta_{h,s}(t))(|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k})}{N-|D|_{k}}\right)^{\delta_{1}}\\ \\ &&\left(1-\displaystyle\frac{(\zeta_{ia,s}(t)+\zeta_{is,s}(t)+\zeta_{h,s}(t))(|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k})}{N-|D|_{k}}\right)^{|S|_{k}-\delta_{1}},\end{array}$ (3.28)
@@
    $\mathbb{P}(I_{a}|E,t)=\begin{pmatrix}|E|_{k}\\ ~\\ \delta_{2}\end{pmatrix}\tau_{e,ia}^{\delta_{2}}(t)(1-\tau_{e,ia}(t))^{|E|_{k}-\delta_{2}},$ (3.30)
@@
    $\mathbb{P}(I_{s}\mid I_{a},t)=\frac{\displaystyle\sum_{\delta_{4}=0}^{|I_{a}|_{k}-\delta_{3}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}(t)^{\delta_{3}}\tau_{ia,r}(t)^{\delta_{4}}(1-\tau_{ia,is}(t)-\tau_{ia,r}(t))^{Q_{1}}}{\displaystyle\sum_{\delta_{3}=0}^{|I_{a}|_{k}}\sum_{\delta_{4}=0}^{|I_{a}|_{k}-\delta_{3}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}(t)^{\delta_{3}}\tau_{ia,r}(t)^{\delta_{4}}(1-\tau_{ia,is}(t)-\tau_{ia,r}(t))^{Q_{1}}},$ (3.31)
@@
    $\mathbb{P}(R\mid I_{a},t)=\frac{\displaystyle\sum_{\delta_{3}=0}^{|I_{a}|_{k}-\delta_{4}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}(t)^{\delta_{3}}\tau_{ia,r}(t)^{\delta_{4}}(1-\tau_{ia,is}(t)-\tau_{ia,r}(t))^{Q_{1}}}{\displaystyle\sum_{\delta_{4}=0}^{|I_{a}|_{k}}\sum_{\delta_{3}=0}^{|I_{a}|_{k}-\delta_{4}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}(t)^{\delta_{3}}\tau_{ia,r}(t)^{\delta_{4}}(1-\tau_{ia,is}(t)-\tau_{ia,r}(t))^{Q_{1}}},$ (3.33)
@@
    $\mathbb{P}(R|I_{s},t)=\frac{\displaystyle\sum_{\delta_{6}=0}^{|I_{s}|_{k}-\delta_{5}}\sum_{\delta_{7}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{6}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}(t)^{\delta_{5}}\tau_{is,h}(t)^{\delta_{6}}\tau_{is,d}(t)^{\delta_{7}}\beta(t)^{Q_{2}}}{\displaystyle\sum_{\delta_{5}=0}^{|I_{s}|_{k}}\sum_{\delta_{6}=0}^{|I_{s}|_{k}-\delta_{5}}\sum_{\delta_{7}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{6}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}(t)^{\delta_{5}}\tau_{is,h}(t)^{\delta_{6}}\tau_{is,d}(t)^{\delta_{7}}\beta(t)^{Q_{2}}}$ (3.34)
@@
    $\mathbb{P}(H|I_{s},t)=\frac{\displaystyle\sum_{\delta_{5}=0}^{|I_{s}|_{k}-\delta_{6}}\sum_{\delta_{7}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{6}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}(t)^{\delta_{5}}\tau_{is,h}(t)^{\delta_{6}}\tau_{is,d}(t)^{\delta_{7}}\beta(t)^{Q_{2}}}{\displaystyle\sum_{\delta_{6}=0}^{|I_{s}|_{k}}\sum_{\delta_{5}=0}^{|I_{s}|_{k}-\delta_{6}}\sum_{\delta_{7}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{6}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}(t)^{\delta_{5}}\tau_{is,h}(t)^{\delta_{6}}\tau_{is,d}(t)^{\delta_{7}}\beta(t)^{Q_{2}}}$ (3.35)
@@
    $\mathbb{P}(D|I_{s},t)=\frac{\displaystyle\sum_{\delta_{5}=0}^{|I_{s}|_{k}-\delta_{7}}\sum_{\delta_{6}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{7}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}(t)^{\delta_{5}}\tau_{is,h}(t)^{\delta_{6}}\tau_{is,d}(t)^{\delta_{7}}\beta(t)^{Q_{2}}}{\displaystyle\sum_{\delta_{7}=0}^{|I_{s}|_{k}}\sum_{\delta_{5}=0}^{|I_{s}|_{k}-\delta_{7}}\sum_{\delta_{6}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{7}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}(t)^{\delta_{5}}\tau_{is,h}(t)^{\delta_{6}}\tau_{is,d}(t)^{\delta_{7}}\beta(t)^{Q_{2}}}$ (3.36)
@@
    $\mathbb{P}(D|H,t)=\frac{\displaystyle\sum_{\delta_{9}=0}^{|H|_{k}-\delta_{8}}\binom{|H|_{k}}{\delta_{8},\delta_{9},|H|_{k}-\delta_{8}-\delta_{9}}\tau_{h,r}(t)^{\delta_{9}}\tau_{h,d}(t)^{\delta_{8}}(1-\tau_{h,r}(t)-\tau_{h,d}(t))^{|H|_{k}-\delta_{8}-\delta_{9}}}{\displaystyle\sum_{\delta_{8}=0}^{|H|_{k}}\sum_{\delta_{9}=0}^{|H|_{k}-\delta_{8}}\binom{|H|_{k}}{\delta_{8},\delta_{9},|H|_{k}-\delta_{8}-\delta_{9}}\tau_{h,r}(t)^{\delta_{9}}\tau_{h,d}(t)^{\delta_{8}}(1-\tau_{h,r}(t)-\tau_{h,d}(t))^{|H|_{k}-\delta_{8}-\delta_{9}}}$ (3.37)
@@
    $\mathbb{P}(R|H,t)=\frac{\displaystyle\sum_{\delta_{8}=0}^{|H|_{k}-\delta_{9}}\binom{|H|_{k}}{\delta_{8},\delta_{9},|H|_{k}-\delta_{8}-\delta_{9}}\tau_{h,r}(t)^{\delta_{9}}\tau_{h,d}(t)^{\delta_{8}}(1-\tau_{h,r}(t)-\tau_{h,d}(t))^{|H|_{k}-\delta_{8}-\delta_{9}}}{\displaystyle\sum_{\delta_{9}=0}^{|H|_{k}}\sum_{\delta_{8}=0}^{|H|_{k}-\delta_{9}}\binom{|H|_{k}}{\delta_{8},\delta_{9},|H|_{k}-\delta_{8}-\delta_{9}}\tau_{h,r}(t)^{\delta_{9}}\tau_{h,d}(t)^{\delta_{8}}(1-\tau_{h,r}(t)-\tau_{h,d}(t))^{|H|_{k}-\delta_{8}-\delta_{9}}}$ (3.39)
@@ 4 COVID-19 Dynamics Over 365 Days
    Figure 3: Low Values of $t$ (from 0 up to 2.182.182.18)
@@
    Figure 4: Moderate to High Values of $t$ (2.182.182.18 and above)
    When $t$ reaches 2.182.182.18 and beyond, the dynamics settle. The susceptible population remains consistently larger than the deceased population, due to a slower increase and decrease in the deceased and susceptible compartments, respectively. This phase signifies a balance between disease spread and recovery, resulting in a more controlled epidemic progression. By approximately day 250, the system reaches a steady state, with variations in all compartments approaching zero. Both the susceptible and deceased populations balance, indicating that the epidemic has subsided and the populations are no longer undergoing significant changes.
These plots illustrate the progression of individuals across different compartments of a COVID-19 model over time (in days), offering valuable insights for understanding and predicting disease dynamics, and for planning public health interventions. Overall, as the probabilistic function increases (when $t$ is in the range [0,0.4)[0,0.4)[ 0 , 0.4 )), there is a rapid decline in the susceptible population, reflecting a high rate of new infections, while the deceased population rises quickly, indicating a severe impact of the disease and a high transition rate from infection to death. This rapid decline in susceptibles and fast increase in deceased population reduce the number of active infections and limit the size of the recovered population, thereby influencing the overall dynamics. When the probabilistic function decreases (for $t$ in the range [0.4,4.2)[0.4,4.2)[ 0.4 , 4.2 )), the decline in the susceptible population slows, indicating a reduction in the infection rate, and the increase in the deceased population also slows, reflecting a less aggressive disease progression. This leads to a more balanced dynamic, with intermediate compartments like the infected population maintaining a more consistent size. Finally, when the probabilistic function becomes constant (approaching 0.150.150.15 for $t$ in the range [4.2,)[4.2$\infty$)[ 4.2 , ∞ )), the sizes of the susceptible and deceased populations settle, indicating that the system has reached an equilibrium with no significant new infections or deaths occurring. This constancy suggests that other compartments, such as the infected and recovered populations, have also leveled off, leading to a long-term steady state in the epidemic’s dynamics.